HETAV PATEL

Grade 11,

Sir what is radiation and what is quantum modelAnd how they emitted

Sir what is radiation and what is quantum modelAnd how they emitted

Grade:11

1 Answers

Arun
6 years ago

Quantum Theory of Radiation

The distribution of energy in the spectrum of radiations of a hot body cannot be explained by applying the classical concepts of physics. Max Planck gave an explanation to this observation by his Quantum Theory of Radiation. His theory says: –

a)      The Radiant energy is always in the form of tiny bundles of light called quanta i.e. the energy is absorbed or emitted discontinuously.

b)      Each quantum has some definite energy E, which depends upon the frequency of the radiations.

E=hγ.

Here, E is the energy of each quantum in Joules, γ is the frequency of the radiations in s-1, h is known as Planck’s constant (a fundamental constant), h=6.626X10-34 J s.

Also, E=hcw, where w is known as wave number. W= (1/λ) m-1.

From above two equations, it is evident that γ=c/λ=cw.

c)       The energy emitted or absorbed by a body is always a whole multiple of a quantum i.e. a body cannot absorb or emit energy in fractions of quantum. This concept is known as quantization of energy.

Edγ= ((8πhγ3)/c3 ) X (dγ)/(exp(hγ/kT)-1)